# Asymptotic regularity of a generalised stochastic Halpern scheme with applications

### v1

@@ 4.1. General results on linear rates
for all $n\in\mathbb{N}$ where $\tilde{A}_{j}^{k}:=\prod_{i=j}^{k}(1-\alpha_{i})$, with $\tilde{A}_{j}^{k}:=1$ for $j>k$. If we furthermore define $\alpha_{n}:=2/(n+2)$ and assume that $s_{0}\leq L$, then
$s_{n}\leq\frac{2L}{n+2}$
for all $n\in\mathbb{N}$.
@@
From the fact that $X_{n+1}\leq(1-\alpha_{n+1})X_{n}+C_{n}$ holds almost surely, we immediately derive
$\mathbb{E}[X_{n+1}]\leq(1-\alpha_{n+1})\mathbb{E}[X_{n}]+\mathbb{E}[C_{n}]$
and Lemma 4.1 yields $\mathbb{E}[X_{n}]\leq 2L/(n+2)$. Proceeding as in the proof of Lemma 2.3, noting that we in particular have $X_{n+1}\leq X_{n}+C_{n}$ almost surely, we similarly derive
@@
so that
$\mathbb{P}(\exists n\geq N(U_{n}\geq\varepsilon))\leq\frac{1}{\varepsilon}\frac{4L}{N+2}.$
This gives
@@ 4.2. Linear rates of asymptotic regularity
Reformulated, the above results in particular state that if $(x_{n}),(y_{n})$ are the sequences generated by (1) for parameters as in (4.2) under the assumption (Hyp) and $\mathbb{E}(\left\lVert\xi_{n}\right\rVert)=O(1/n^{2})$ as well as $\mathbb{E}(\left\lVert\delta_{n}\right\rVert)=O(1/n^{2})$, then
$\mathbb{E}[\left\lVert x_{n}-x_{n+1}\right\rVert]=O(1/n)\text{ and }\mathbb{E}[\left\lVert y_{n}-y_{n+1}\right\rVert]=O(1/n).$
@@ 4.3. Linear rates of asymptotic regularity relative to the mappings in special cases
In the special case of the stochastic Halpern iteration, which we reobtain (as discussed before) by setting $U:=\mathrm{Id}$ as well as $\delta_{n}:=0$, we get the following fast rates in the above special case:
@@
Theorem 4.8 is closely related to [4, Theorem 3.3], but with adjusted step-sizes that now provide exact linear rates (without logarithmic factors).
@@ 4.4. Fast rates of asymptotic regularity relative to the mappings in the general case
First note that in the context of moduli $\eta$ of power type $p\geq 2$, we are actually in the setting of the previous Remarks 3.13 and 3.19 where $\tilde{\eta}(\varepsilon)=C\varepsilon^{p-1}$. Then the rate for $\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$ almost surely follows by instantiating the rate given in Theorem 3.12 with the following moduli: With the above $\tilde{\eta}$, we have $\widehat{\varepsilon}:=C\Lambda^{2}\varepsilon^{p}\lambda^{p-1}/(3K)^{p-1}$ and $K^{\prime}:=3K/\lambda$. Using Theorem 4.4, we have
$\mathbb{P}\left(\exists i\geq n\left(\left\lVert x_{i}-x_{i+1}\right\rVert\geq\varepsilon\right)\right)\leq\frac{1}{\varepsilon}\frac{4L_{0}}{n+2}$
for all $n\in\mathbb{N}$ and a suitable constant $L_{0}$ arising as an integer linear combination of $K$, $K_{1}$ and $K_{2}$. So we in particular have that $\Delta(\lambda,\varepsilon)=\left\lceil 4L_{0}/\varepsilon\lambda\right\rceil$ is a corresponding rate for $\left\lVert x_{n}-x_{n+1}\right\rVert\to 0$ almost surely. As $\alpha=2/(n+2)$, we further have rather immediately that $\rho(\varepsilon)=\left\lceil 2/\varepsilon\right\rceil$ is a corresponding rate for $\alpha_{n}\to 0$. Lastly, using the assumptions on $\left\lVert\xi_{n}\right\rVert$ and $\left\lVert\delta_{n}\right\rVert$, note that as in Remark 4.5 we have $\sum_{n=N}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert]\leq K_{1}/(N+1)$ so that $\chi_{1}(\varepsilon)=\left\lceil K_{1}/\varepsilon\right\rceil$ is a corresponding rate of convergence for $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert]<\infty$. The rate of convergence $\chi_{2}(\varepsilon)=\left\lceil K_{2}/\varepsilon\right\rceil$ for $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\delta_{n}\right\rVert]<\infty$ follows similarly. Then instantiating Theorem 3.12 under Remark 3.13 gives us the rate
@@
5. Further remarks
We conclude the paper with an extended discussion on two further topics: oracle complexities and applications in reinforcement learning. While both of these topics are being explored in greater depth in ongoing future work, we want to sketch key aspects of these developments here.
5.1. Abstract oracle complexities
The intuition behind our abstract stochastic scheme is that $\xi_{n}$ and $\delta_{n}$ arise by evaluating stochastic oracles $\tilde{T}$ and $\tilde{U}$ for $T$ and $U$ respectively. Though our primary focus has been on establishing direct convergence rates for our schemes, in concrete applications the resulting oracle complexity may act as a more reliable estimate of the actual cost of running the algorithm. To address this at a level of generality in line with our overall approach, we propose an abstract characterisation of oracle complexity. We first introducing a pair of sequences of natural numbers (#$\#\xi_{n}$) and (#$\#\delta_{n}$), with the intuition that #$\#\xi_{n}$ denotes the number of calls made to the stochastic oracle $\tilde{T}$ in computing the approximation to $Tx_{n}$ (with output $Tx_{n}+\xi_{n}$), and #$\#\delta_{n}$ the number of calls to $\tilde{U}$ in computing the approximation to $Uy_{n}$.
With this intuition in mind, we say that $\psi:\mathbb{N}\to\mathbb{N}$ is a bound on the oracle growth of the general scheme (1) if
$\forall N\in\mathbb{N}\,\left(\#\xi_{N}+\sum_{n=0}^{N-1}(\#\xi_{n}+\#\delta_{n})\leq\psi(N)\right)$
so that the total number of oracle calls needed to compute $x_{N}$ and $y_{N}$ is bounded by $\psi(N)$. Then, for example, if
$\mathbb{E}[\left\lVert Tx_{n}-x_{n}\right\rVert]\to 0$ with rate $\varphi$,
we can compute for any $\varepsilon>0$ an approximant $x_{N}$ such that
$\mathbb{E}[\left\lVert Tx_{N}-x_{N}\right\rVert]<\varepsilon$ using at most $(\psi\circ\varphi)(\varepsilon)$ queries to the stochastic oracles.
Indeed, the bounding function $\psi$ can be used in a similar way to convert any of our convergence results to a corresponding characterisation of the overall oracle complexity in a completely general manner. One possible instantiation of this abstract perspective is discussed in the following example:
Example 5.1.
Assume that $\tilde{T}$ is evaluated via a minibatch strategy (as e.g. studied in [4]), which would correspond to setting
$\xi_{n}:=\frac{1}{k_{n}}\sum_{j=1}^{k_{n}}\tilde{T}(x_{n},\zeta_{n,j})-Tx_{n}$
for batchsizes $(k_{n})$ (where $\zeta_{n,1},\ldots,\zeta_{n,k_{n}}$ are i.i.d. random variables sampled from a distribution associated with $x_{n}$) and therefore #$\#\xi_{n}=k_{n}$. Under suitable additional variance assumptions on the oracle $\tilde{T}$, one can then show that $\mathbb{E}[\left\lVert\xi_{n}\right\rVert]\leq c/\sqrt{k_{n}}$ for some corresponding constant $c>0$. Hence, in such situations, the oracle bounds then involve a payoff between the rate of variance reduction and the oracle growth, or more concretely between the
($\ast$) convergence speed of 1/kn0 and growth of $\text{convergence speed of $1/\sqrt{k_{n}}\to 0$ and growth of $\sum_{n=0}^{N}k_{n}$}.$
One concrete such situation is e.g. discussed in [4, Corollary 3.5] where, in the context of a finite dimensional space $X$, the oracle $\tilde{T}$ is assumed to have uniformly bounded variance in the sense that
$\sup_{x\in X}\mathbb{E}\left[\left\lVert\tilde{T}(x,\zeta)-Tx\right\rVert^{2}_{2}\right]\leq\sigma^{2}$
[…]
In the special cases of the stochastic Halpern iteration ($U:=\mathrm{Id}$ and $\delta_{n}:=0$) or the stochastic Krasnoselskii-Mann iteration with Tikhonov regularization terms ($T:=\mathrm{Id}$ and $\xi_{n}:=0$), under the additional conditions (Hyp) with constant $K_{0}$ on the boundedness and (4.2) on the parameters as in Theorems 4.8 and 4.9, said theorems immediately imply
$\mathbb{E}[\left\lVert Tx_{n}-x_{n}\right\rVert]\leq\frac{2L}{n+2}\text{ and }\mathbb{E}[\left\lVert Ux_{n}-x_{n}\right\rVert]\leq\frac{2L^{\prime}}{n+2}$
for suitable $L$ and $L^{\prime}$ constructed from $K_{0},K_{1}$ or $K_{0},K_{2}$ and $B\geq 1/(1-\beta)$, respectively. These estimates correspond to rates of convergence $\varphi(\varepsilon):=\varepsilon/2L$ and $\varphi^{\prime}(\varepsilon):=\varepsilon/2L^{\prime}$, respectively, and therefore, in that context, we can compute an $x_{N}$ such that
$\mathbb{E}[\left\lVert Tx_{N}-x_{N}\right\rVert]<\varepsilon\text{ and }\mathbb{E}[\left\lVert Ux_{N}-x_{N}\right\rVert]<\varepsilon$
using at most $\left(\varepsilon/2L\right)^{5}$ or respectively $\left(\varepsilon/2L^{\prime}\right)^{5}$ queries to the stochastic oracles $\tilde{T}$ or $\tilde{U}$ in question. In this way, we in particular obtain a version of [4, Corollary 3.5] without logarithmic factors.
As is to be expected when augmenting existing deterministic methods with noise terms, all our convergence results require us to control the means of $(\xi_{n})$ and $(\delta_{n})$ in some way, and this is intuitively achieved in concrete settings by increasing the number of oracle queries (#$(\#\xi_{n})$ and (#$(\#\delta_{n})$, so that, as above with ($\ast$5.1), obtaining good bounds on the oracle complexity for instances of (1) will generally involve a tension between variance control and oracle growth. Because our explicit rates of convergence make completely precise how improved control on the means through e.g. faster convergence rates for $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert]$ and $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\delta_{n}\right\rVert]$ leads to improved rates of convergence for the algorithm itself, we anticipate that in any given model that allows us to explicitly describe $\psi$ in terms of $\mathbb{E}[\left\lVert\xi_{n}\right\rVert]$ and $\mathbb{E}[\left\lVert\delta_{n}\right\rVert]$, our abstract quantitative results would not only provide us with a general bound on the associated oracle complexity in terms of parameters representing variance control, but might even provide insights into specific choices of parameters that optimize oracle complexity.
5.2. Applications in reinforcement learning
Our new stochastic iteration scheme (1) immediately leads to new algorithms in relevant areas of application, including reinforcement learning, where in particular noisy methods for computing fixpoints underlie model-free learning algorithms such as $Q$-learning [39]. Here, direct access to the corresponding operators $T,U$ would require full knowledge of the agent’s environment, so these are instead sampled, leading to iteration schemes with noise terms. Only very recently in [4] has the special case of the scheme (1) corresponding to Halpern’s iteration (obtained by setting $U:=\text{Id}$ and $\delta_{n}:=0$ in (1)) been instantiated to yield a Halpern-type version of $Q$-learning, and in a similar way our general method represents an expanded class of learning algorithms.
@@ 5.2. Applications in reinforcement learning
To illustrate this, suppose that $(S,A,r,p)$ forms a Markov decision process (MDP) over some finite set of states $S$ and actions $A$, where if we choose action $a$ in state $s$, $r(s,a)$ represents an immediate reward and $p(s,a,t)$ the probability that we transition to state $t$. Concretely, by appropriately instantiating our stochastic Krasnoselskii-Mann iteration with Tikhonov regularization terms (obtained by setting $T:=\text{Id}$ and $\xi_{n}:=0$ as well as $\gamma_{n}:=1-\alpha_{n}$ and $u:=0$ in (1)), we are lead to the following novel $Q$-learning method with Tikhonov regularization terms:
(KM-T-Q) $Q_{n+1}(s,a):=(1-\beta_{n})\left(U\left(\gamma_{n}Q_{n}\right)(s,a)+\delta_{n}(s,a)\right)+\beta_{n}\left(\gamma_{n}Q_{n}(s,a)\right)$
for
$(UQ)(s,a):=r(s,a)+\sum_{t\in S}p(s,a,t)\max_{b\in A}Q(t,b)+c$
for some constant $c\in\mathbb{R}$, where here $U$ is clearly nonexpansive with respect to
$\left\lVert Q\right\rVert_{\infty}=\max_{s\in\mathcal{S}}\max_{a\in\mathcal{A}}Q(s,a).$
There are standard conditions on our MDP (see e.g. [35]) under which an optimal average reward, i.e. a maximum $v^{\ast}$ for
$v_{\pi}:=\liminf_{N\to\infty}\mathbb{E}\left[\frac{1}{N}\sum_{n=0}^{N-1}r(s_{n},a_{n})\right]\text{with $(s_{n},a_{n})$ induced by the policy $\pi$}$
exists and is independent of the initial state, and moreover, an optimal policy obtaining this maximum can be defined as
$\pi^{\ast}(s):=\operatorname*{arg\,max}_{a\in A}Q^{\ast}(s,a)$
where $Q^{\ast}$ is a fixpoint of $U$ for $c:=-v^{\ast}$. Thus the convergence of (KM-T-Q) and similar schemes towards fixpoints is of direct practical relevance in computing optimal policies. In particular, instantiating the noise terms according to a regularized minibatch strategy as e.g. discussed in [4] for a Halpern-type version of $Q$-learning, i.e.
$\delta_{n}(s,a):=\gamma_{n}\left(\frac{1}{s_{n}}\sum_{j=1}^{s_{n}}\max_{b\in S}Q_{n}(\zeta_{n,j},b)-\sum_{t\in S}p(s,a,t)\max_{b\in A}Q_{n}(t,b)\right)$
@@
where $\zeta_{n,1}\ldots,\zeta_{n,k_{n}}$ are i.i.d. random variables sampled from $p(s,a,\cdot)$, (KM-T-Q) then becomes
$Q_{n+1}(s,a):=\gamma_{n}\left((1-\beta_{n})\left(r(s,a)+\frac{1}{s_{n}}\sum_{j=1}^{s_{n}}\max_{b\in S}Q_{n}(\zeta_{n,j},b)-v^{\ast}\right)+\beta_{n}Q_{n}(s,a)\right)$
the convergence of which can then be used (as in [4]) to establish the convergence of concrete variants of RVI-$Q$-learning in the style of the Krasnoselskii-Mann iteration with Tikhonov regularization terms, where in the above iteration $v^{\ast}$ is replaced by some approximant $f(Q_{n})$ for a suitable function $f$.
However, our focus in this paper is not the development of concrete learnings algorithms, but the establishment of abstract convergence results that are of general relevance across areas of application. In that spirit, we simply observe for now that convergence of the abstract $Q$-learning scheme (KM-T-Q) follows immediately from our main convergence results and inherits the associated rates. For example, in the case of fast rates we obtain the following:
Corollary 5.2.
Let $(Q_{n})$ be the sequence generated by (KM-T-Q) for the case of parameters
$\gamma_{n}=1-\frac{2}{n+2}\text{ and }\beta_{n}=\beta\in(0,1).$
Assume that $\left\lVert Q_{n}\right\rVert_{\infty},\left\lVert r+c\right\rVert_{\infty}\leq K$ so that (Hyp) holds with constant $K_{0}:=2K$. Also, assume that $\mathbb{E}[\left\lVert\delta_{n}\right\rVert_{\infty}]\leq K_{2}/(n+2)^{2}$. Lastly, let $B\geq 1/(1-\beta)$. Then
$\mathbb{E}[\left\lVert UQ_{n}-Q_{n}\right\rVert_{\infty}]\leq\frac{2L}{n+2}$
for all $n\in\mathbb{N}^{*}$ where $L$ is constructed from $K_{0},K_{2}$ and $B$ as in Theorem 4.9.
Remark 5.3.
Instantiating $\delta_{n}$ as above via a regularized minibatch strategy, under conditions as described in Example 5.1 we would obtain the desired control of the means by defining our batchsizes through $s_{n}:=(n+4)^{4}$, which akin to the discussion in Example 5.1 also immediately would allow us an estimation of the oracle complexity in that case which is of pentic complexity.
It would be very interesting to see in future work whether $Q$-learning could, by means of Tikhonov regularization terms as in the scheme (KM-T-Q), benefit from the nice theoretical and practical properties that these regularization terms induce in the non-stochastic case.
We conclude by remarking that much broader generalisations which more fully exploit the convergence results of this paper are possible. One obvious avenue for further study would be to consider alternating $Q$-learning algorithms based on our main scheme (1) and we want to propose on such example by considering the following scheme which represents a Halpern-Mann-type variant of $Q$-learning (where the anchor point is taken to be zero):
(sHM-Q) $Q^{\prime}_{n}(s,a):=(1-\alpha_{n})\left(r(s,a)+\sum_{t\in S}p(s,a,t)\max_{b\in A}Q_{n}(t,b)+\xi_{n}(s,a)\right)$
$Q_{n+1}(s,a):=(1-\beta_{n})\left(r^{\prime}(s,a)+\sum_{t\in S}p^{\prime}(s,a,t)\max_{b\in A}Q^{\prime}_{n}(t,b)+\delta_{n}(s,a)\right)+\beta_{n}Q^{\prime}_{n}$
where $(S,A,r^{\prime},p^{\prime})$ is a second MDP over the same states and actions. This method bears a passing resemblance to double $Q$-learning [16], which incorporates a double estimator to reduce bias in ordinary $Q$-learning. Moreover, our scheme converges to a simultaneous fixpoint of two underlying nonexpansive operators, and thus might be relevant in situations where we are required to compute optimal policies concurrently across distinct environments. However, we leave an exploration of the potential merits of both schemes (KM-T-Q) and (sHM-Q) as well as similar algorithms to future work.
Acknowledgments: The first author was supported by the ‘Deutsche Forschungsgemeinschaft’ Project DFG KO 1737/6-2. The second author was supported by the EPSRC grant EP/W035847/1.

### v2

@@ 4.1. General results on linear rates
for all $n\in\mathbb{N}$ where $\tilde{A}_{j}^{k}:=\prod_{i=j}^{k}(1-\alpha_{i})$, with $\tilde{A}_{j}^{k}:=1$ for $j>k$. If we furthermore define $\alpha_{n}:=2/(n+2)$ and assume that $s_{0}\leq L$, then $s_{n}\leq 2L/(n+2)$ for all $n\in\mathbb{N}$.
@@
From the fact that $X_{n+1}\leq(1-\alpha_{n+1})X_{n}+C_{n}$ holds almost surely, we immediately derive $\mathbb{E}[X_{n+1}]\leq(1-\alpha_{n+1})\mathbb{E}[X_{n}]+\mathbb{E}[C_{n}]$ and Lemma 4.1 yields $\mathbb{E}[X_{n}]\leq 2L/(n+2)$. Proceeding as in the proof of Lemma 2.3, noting that we in particular have $X_{n+1}\leq X_{n}+C_{n}$ almost surely, we similarly derive
@@
so that $\mathbb{P}(\exists n\geq N(U_{n}\geq\varepsilon))\leq\frac{1}{\varepsilon}\frac{4L}{N+2}$. This gives
@@ 4.2. Linear rates of asymptotic regularity
Reformulated, the above results in particular state that if $(x_{n}),(y_{n})$ are the sequences generated by (1) for parameters as in (4.2) under the assumption (Hyp) and $\mathbb{E}(\left\lVert\xi_{n}\right\rVert)=O(1/n^{2})$ as well as $\mathbb{E}(\left\lVert\delta_{n}\right\rVert)=O(1/n^{2})$, then $\mathbb{E}[\left\lVert x_{n}-x_{n+1}\right\rVert]=O(1/n)$ as well as $\mathbb{E}[\left\lVert y_{n}-y_{n+1}\right\rVert]=O(1/n)$.
@@ 4.3. Linear rates of asymptotic regularity relative to the mappings in special cases
In the special case of the stochastic Halpern iteration, which we reobtain (as discussed before) by setting $U:=\mathrm{Id}$ as well as $\delta_{n}:=0$, we get the following fast rates:
@@
Theorem 4.8 is closely related to [5, Theorem 3.3], but with adjusted step-sizes that now provide exact linear rates (without logarithmic factors).
@@ 4.4. Fast rates of asymptotic regularity relative to the mappings in the general case
First note that in the context of moduli $\eta$ of power type $p\geq 2$, we are actually in the setting of the previous Remarks 3.13 and 3.19 where $\tilde{\eta}(\varepsilon)=C\varepsilon^{p-1}$. Then the rate for $\left\lVert Uy_{n}-y_{n}\right\rVert\to 0$ almost surely follows by instantiating the rate given in Theorem 3.12 with the following moduli: With the above $\tilde{\eta}$, we have $\widehat{\varepsilon}:=C\Lambda^{2}\varepsilon^{p}\lambda^{p-1}/(3K)^{p-1}$ and $K^{\prime}:=3K/\lambda$. Using Theorem 4.4, we have $\mathbb{P}\left(\exists i\geq n\left(\left\lVert x_{i}-x_{i+1}\right\rVert\geq\varepsilon\right)\right)\leq\frac{1}{\varepsilon}\frac{4L_{0}}{n+2}$ for all $n\in\mathbb{N}$ and a suitable constant $L_{0}$ arising as an integer linear combination of $K$, $K_{1}$ and $K_{2}$. So we in particular have that $\Delta(\lambda,\varepsilon)=\left\lceil 4L_{0}/\varepsilon\lambda\right\rceil$ is a corresponding rate for $\left\lVert x_{n}-x_{n+1}\right\rVert\to 0$ almost surely. As $\alpha=2/(n+2)$, we further have rather immediately that $\rho(\varepsilon)=\left\lceil 2/\varepsilon\right\rceil$ is a corresponding rate for $\alpha_{n}\to 0$. Lastly, using the assumptions on $\left\lVert\xi_{n}\right\rVert$ and $\left\lVert\delta_{n}\right\rVert$, note that as in Remark 4.5 we have $\sum_{n=N}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert]\leq K_{1}/(N+1)$ so that $\chi_{1}(\varepsilon)=\left\lceil K_{1}/\varepsilon\right\rceil$ is a corresponding rate of convergence for $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\xi_{n}\right\rVert]<\infty$. The rate of convergence $\chi_{2}(\varepsilon)=\left\lceil K_{2}/\varepsilon\right\rceil$ for $\sum_{n=0}^{\infty}\mathbb{E}[\left\lVert\delta_{n}\right\rVert]<\infty$ follows similarly. Then instantiating Theorem 3.12 under Remark 3.13 gives us the rate
@@
5. An outlook onto future work
We conclude the paper with a brief discussion on the potential of the schemas and results given here and, with that, an outlook onto forthcoming work. Concretely, in a forthcoming paper, we will outline applications of the present paper to problems in reinforcement learning, where a novel variant of the prominent model-free $Q$-learning algorithm [41] with Tikhonov regularization terms is presented, and the asymptotic regularity results from the main theoretical part of this paper are utilized to give very good oracle complexity estimates for a specific mini-batching procedure that matches those recently obtained for Halpern’s iteration in [5]. In more detail, only very recently in [5] has the special case of the scheme (1) corresponding to Halpern’s iteration (obtained by setting $U:=\text{Id}$ and $\delta_{n}:=0$ in (1)) been instantiated to yield a Halpern-type version of $Q$-learning, and in a similar way our general method represents an expanded class of learning algorithms. A key aspect of those variants, already motivating [5], is that they stay computationally effective over the averaged reward setting as the iterations allow for fast asymptotic behavior even in the presence of general nonexpansive maps.
@@ 5. An outlook onto future work
In the case of our stochastic Krasnoselskii-Mann iteration with Tikhonov regularization terms (obtained by setting $T:=\text{Id}$ and $\xi_{n}:=0$ as well as $\gamma_{n}:=1-\alpha_{n}$ and $u:=0$ in (1)), this in particular leads to the following novel $Q$-learning method with Tikhonov regularization terms: Let $(S,A,r,p)$ be a given Markov decision process with a finite set of states $S$ and actions $A$, where if we choose action $a$ in state $s$, $r(s,a)$ represents an immediate reward and $p(s,a,t)$ the probability that we transition to state $t$. Under standard conditions on the Markov decision process and by instantiating the noise terms according to a regularized minibatch strategy as e.g. discussed in [5] for a Halpern-type version of $Q$-learning, we can then utilize variants of RVI-$Q$-learning [1] in the style of the Krasnoselskii-Mann iteration with Tikhonov regularization terms to produce the new scheme
$Q_{n+1}(s,a):=(1-\beta_{n})\left(r(s,a)+\frac{\gamma_{n}}{s_{n}}\sum_{j=1}^{s_{n}}\max_{b\in S}Q_{n}(\zeta_{n,j},b)-f(\gamma_{n}Q_{n}(s,a))\right)+\beta_{n}\gamma_{n}Q_{n}(s,a),$
@@
to quickly approximate the optimal policy in the average reward setting, where $\zeta_{n,1}\ldots,\zeta_{n,s_{n}}$ are i.i.d. random variables sampled from $p(s,a,\cdot)$, and $f$ is a suitable function (see e.g. [1]).
A detailed exposition of these new methods, and of the use of the additional degree of freedom gained by the Tikhonov regularization terms, together with a careful and rather general examination of associated oracle complexities, will be given in a forthcoming paper. In that paper, we will also examine a novel use of the second mapping in the schema (1) in the style of double Q-learning [17], optimizing over two concurrent Markov decision processes in parallel environments.